Given your $1,000 investment at 5% simple interest, how much money will you have after 2 years?
Answer #1

Answer

We can calculate the amount with following formula

A=p (1+rt)

A=1000(1+(0.05*2)

A=1000(1.10)

A=1100

Explanation

Here, A=total amount after 2 year

P=amount you invest

R=interest rate

T=time (number of years)